OpenStudy (marziajackson$):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2 h, and Car B traveled the distance in 1.5 h. Car B traveled 15 mph faster than Car A. How fast did Car B travel?

OpenStudy (anonymous):

45mph

OpenStudy (anonymous):

just make a proportion

OpenStudy (anonymous):

\[\frac{ 2 }{ x-15 }=\frac{ 1.5 }{ x }\]

OpenStudy (anonymous):

and solve

OpenStudy (anonymous):

cool? @MarziaJackson$

OpenStudy (marziajackson$):

@Alicia_redd